find the general solution of the equation tan(mx)+cot(nx)=0?
 

Grade:11

1 Answers

sagar kumar
6 years ago
tanmx+cotnx=0sinmx.sinnx+cosmx.cosnx=0cos(m-n)x=0(m-n)x=(2k-1/2)n,k belongs to integerx=[2k+1/2(m-n)] ,here k belongs to integer